# The Serial Scaling Hypothesis

### v2

@@ 4 Diffusion model’s computation is not serial
The Serial Scaling Hypothesis provide a lens to look at machine learning from serial–parallel lens. We have examined serial machine learning problems. We now turn to a diffusion models (Ho et al., 2020; Song et al., 2021; Song & Ermon, 2019), widely used in image/video generation (Brooks et al., 2024) and other vision tasks such as depth estimation (Saxena et al., 2023), and language modeling tasks (Li et al., 2022; Nie et al., 2025; Arriola et al., 2025). While commonly thought as a serial model, we show here for the first time that they are not.
@@
While the number of denoising steps $T$ is scalable—more denoising steps for finer approximation of $p_{\text{truth}}$, in practice, diffusion models converge rapidly (Ma et al., 2025). For instance, image generation plateaus at 300 steps (Nichol & Dhariwal, 2021), depth estimation shows little difference between 5 and 100 steps (Ravishankar et al., 2024), and language modeling yields similar perplexity for 32 vs. 1024 steps (Austin et al., 2021). With distillation, 1–4 steps suffice without much loss (Yin et al., 2024; Liu et al., 2024; Song et al., 2023; Lin et al., 2024; Salimans & Ho, 2022). Such rapid convergence suggests that the effective computation depth of diffusion models is low. It would be surprising if the underlying computation were truly serial.
Diffusion models with a $\mathsf{TC}^{0}$ backbone can only solve problems in $\mathsf{TC}^{0}$. Previous work showed that a backward diffusion process converges to $p_{0}$ at the rate $TV=\mathcal{O}(d/T)$ (Li & Yan, 2024, Thm. 1), where $TV$ is the total variation between $p_{0}$ and the denoising $p_{\theta,0}$, and $d$ is the dimension of $x_{N+1}$. Building on top of it, we obtained the following theorem, the formal statement and proof of which are in Appendix F:
@@
What does this mean? Diffusion models only provide a constant amount of additional serial computation. The above theorem precludes the use of diffusion models as a scalable means of increasing serial computation. Unlike CoT, which genuinely adds serial compute (see Appendix˜E), diffusion models do not.
Limitation. The theorem doesn’t apply: (1) If the dimension of $x_{n+1}$ is increasing. However, diffusion language modeling, as output length increases, has yet to show promising results (Austin et al., 2021; Lou et al., 2024; Gulrajani & Hashimoto, 2023; Sahoo et al., 2024). (2) If the backbone is either poorly trained or trained under a different objective. This leaves open the possibility of serial-compute-scalable generative models beyond score-based diffusion.
@@ 5 Acknowledgements
We are grateful to William Merrill and Anant Sahai for their invaluable technical feedback and corrections that significantly strengthened our theoretical analysis. We also thank Angjoo Kanazawa, Justin Kerr, Amil Dravid, Yossi Gandelsman, and Yuatyong Chaichana for their thoughtful comments and suggestions that helped improve the clarity and presentation of this work. KP is supported by the ONR MURI grant.
@@ Appendix C Potential Misconceptions
One may be confused by the statement that “Transformers are not serial,” followed by “Chain of Thought is serial.” This appears to be a contradiction. It is not, though it may appear so due to our omission of words. The first statement really should be “If you fix a Transformer’s parameters and give it problems of size $\mathcal{O}(n)$, but only run it for $\mathcal{O}(1)$ forward passes, then you cannot solve inherently serial problem-families.” Similarly, the second statement really should be “If you fix a Transformer’s parameters and give it problems of size $\mathcal{O}(n)$, and also run it for $\mathcal{O}(n)$ forward passes, then you can solve inherently serial problem-families.”
This is a special case of a general phenomenon: people confuse architecture and inference methods, meshing them all together like two lumps of clay gently melding under the desert sun. This is unfortunate but understandable because, in practice, architecture and inference couple together strongly, and this confusion is also encouraged by things like “architecture–deployment co-design.” The fact is that the abstractions leak. How you use a model for inference depends on its architecture. How you design a new model architecture depends on what you expect its inference method to be. This is a good thing, but it may cause people to confuse architecture and inference.
Fundamentally, we consider an “architecture” to be a class, and its instances are “atoms of computation,” and this is the intuition that the reader should keep in mind. An atom has mass and volume. An atom can be copied and put into many molecules. What an atom cannot do is be divided or changed. Similarly, an instance of the “Transformer architecture” is an “atomic” Transformer. This atomic Transformer does have subatomic structures, such as its individual layers. However, since we are not performing complicated model surgeries, such as taking off its head, reading out the tensor from its middle layers, etc., we can assume it is atomic. The least you can do with this atomic Transformer is perform one single forward pass, and this has a definite computational cost. It may be 48 layers or 96 layers, but the cost is definite.
@@ C.2 SSM is not RNN
We class SSMs as not inherently serial, while RNNs as inherently serial. This is not a mistake, and in fact, a very valuable illustration of how an architecture can appear serial but is not.
@@
So they removed recurrence and preserved attention. Their title meant “attention is all you need (and recurrence is not).” This is exactly what our paper is arguing against: attention is not all you need, and inherently serial processing is necessary. This may be performed by recurrence or some other architecture, but it cannot be avoided.
@@
Concretely, consider the problem of ingesting an input sequence: $x_{1},x_{2},\dots,x_{n}$, and the task is to output the next token $x_{n+1}$. For a Transformer with 96 layers to accomplish the task, it does so by 96 sequential steps, each step is very wide but not deep. In contrast, an RNN with 96 layers must process $x_{1}$ first, then $x_{2}$, and so on. It requires $96n$ sequential steps that cannot be done in parallel. Suppose one attempts to do them “in parallel,” then there is an immediate problem: To run the RNN on $x_{2}$, one needs the internal states of the RNN just after it has processed $x_{3}$, and to do that, it needs to have processed $x_{2}$, etc. Abstractly, the operation of an RNN looks like:
@@
In short, because the internal states of an RNN change nonlinearly and unpredictably, one cannot skip the steps.
@@ Appendix D $\mathsf{TC}^{0}$ and $\mathsf{TC}$ Classes
Appendix E Computation capabilities of modern machine learning methods
In this section, we point out an observational fact that machine learning methods that allow parallelization of their computation graph, such as MLPs, Transformers, and SSMs, have so far been shown to lack the ability to solve inherently serial problems. On the contrary, the only methods that are capable of solving serial problems are those that cannot be parallelized, such as RNNs, repeating layers, and CoT.
@@ Appendix E Computation capabilities of modern machine learning methods
In terms of computational complexity, MLPs have been formalized as a family of Boolean circuits with threshold gates called $\sf{TC}^{0}$ (Parberry & Schnitger, 1988). Consider a family of MLPs with constant $\mathcal{O}(1)$ depth, $\mathsf{poly}(n)$ neurons per layer, and $\mathcal{O}(1)$ numerical precision. A problem is in $\sf{TC}^{0}$ if and only if it is decidable by one forward pass through an MLP under this setup.
@@
It is widely suspected that $\sf{TC}^{1}$ is strictly larger than $\sf{TC}^{0}$. Therefore, any $\sf{TC}^{1}$-hard problem requires computations deepening at a rate of at least $\mathcal{O}(\ln n)$. We may consider the problem as being serial in a weaker, logarithmic sense. A simple example of $\sf{TC}^{1}$ is the word problem of the symmetric group on 5 elements $S_{5}$: Given $g_{1},g_{2},\dots,g_{n}\in S_{5}$, find $\prod_{i}g_{n}$ (Liu et al., 2022). Intuitively, this is because $S_{5}$ is not a solvable group, and therefore there is no better way to multiply its elements than via a binary tree, which has $\mathcal{O}(\ln n)$ layers. Attempting to perform this with just $\mathcal{O}(1)$ layers would effectively require one to memorize the entire $S_{5}^{n}\to S_{n}$ multiplication table, which scales exponentially as $e^{n\ln(120)}$, an exponential depth-width trade-off.
@@
To solve the word problem of $S_{5}$ by a Transformer in one forward pass, we present it with input tokens $g_{1},g_{2},\dots,g_{n},y$, where $y$ is a special token. The Transformer’s output at $y$ is read out as the answer. Similarly, we may solve such a problem by presenting the same input tokens to a State Space Model (SSM) or a Recurrent Neural Network (RNN). Since a Transformer with $\mathcal{O}(1)$ layers and $\mathsf{poly}(n)$ dimensions can only solve problems in $\sf{TC}^{0}$ (Merrill & Sabharwal, 2023b) in one forward pass, it cannot solve the word problem of $S_{5}$ that lies outside $\sf{TC}^{0}$.
@@
On the contrary, for an RNN, we can write the multiplication table of $S_{5}^{2}\to S_{5}$ directly into its weights, so it can solve the problem by unrolling for $\mathcal{O}(n)$ recurrence steps with $\mathcal{O}(1)$ layers and $\mathcal{O}(1)$ dimensions. Intuitively, the hidden states of an RNN keep track of the progress of multiplication as it performs the forward passes. However, RNN’s recurrence state dependency renders it non-parallelizable. Several families of SSMs were developed as a compromise that still have recurrence on hidden states, like an RNN, while making forward passes parallelizable, like a Transformer. The prototypical SSM architecture is Mamba (Gu & Dao, 2023), though there are many others.
@@
Unfortunately, it has been proven that there is not yet a “free serial-compute lunch,” in the sense that the main families of SSMs proposed so far still cannot solve the word problem of $S_{5}$ under the constraints of $\mathcal{O}(1)$ layers, $\mathsf{poly}(n)$ dimensions per layer, and one forward pass. Despite its apparent recurrence, the hidden state offered by an SSM is weaker in a computational sense than that offered by an RNN (Strobl et al., 2023; Merrill & Sabharwal, 2023a; Merrill et al., 2024). This theoretical fact rhymes with the empirical fact that in practice, MLPs, Transformers, and SSMs are all more parallelizable than RNNs.
@@
Only a genuinely serial method has been shown to go beyond the $\mathsf{TC}^{0}$ class. In addition to RNNs, this includes repeating layers and CoT. Repeating layers for $\mathcal{O}(\ln n)$ times enables a standard Transformer to solve tasks in the $\mathsf{TC}^{1}$ class (Merrill & Sabharwal, 2025). With $\mathsf{poly}(n)$ CoT, requiring multiple forward passes before producing a final answer, a Transformer can be lifted from $\mathsf{TC}^{0}$ to $\mathsf{P}$ (Feng et al., 2023; Li et al., 2024). As discussed in Section˜3.5, the power of CoT has been well-attested in practice by the improved performance of reasoning models in complex math and science tasks.
@@
Appendix F Diffusion is in non-uniform $\mathsf{TC}^{0}$
@@
F.1 Language modeling
At the most abstract level, a language is simply a set of words made of letters. Formally:
@@ F.1 Language modeling
  • A word in an alphabet $\Sigma$ is a finite sequence of elements of $\Sigma$.
  • A language $L$ in an alphabet $\Sigma$ is a set of words in the alphabet $\Sigma$.
@@
A prefix language modeling problem is, given a sequence of tokens $x_{1},\dots,x_{n}$, to compute the next token $x_{n+1}$. This is a deterministic formalization of “next-token prediction,” the dominant paradigm in language modeling since the GPT-2 of 2019. This is also a decision problem formulation: input of size $n$, $x_{1},\dots,x_{n}$, and a decision of either “yes” or “no,” $x_{n+1}$.
While usually, a diffusion model is used for generating from a continuous state space such as $\mathbb{R}^{d}$, it can be used to model discrete distributions as well. This is necessary for language modeling. We consider the case closest to continuous state space modeling—quantization: One divides the continuous state space $\mathbb{R}^{d}$ into regions and assigns a token to each region. This then allows sampling a discrete distribution from a diffusion model with continuous state space. Formally, if $\Sigma=\{a_{1},a_{2},\dots,a_{M}\}$ is the alphabet, then we divide $\mathbb{R}^{d}$ into $M$ regions $V_{1},\dots,V_{M}$, such that each region $V_{i}$ maps to a token $a_{i}$.
In non-uniform circuit complexity, for each task instance $n$, we have a $\mathsf{TC}^{0}$ score-network $f_{\theta,n}$ such that:

### v3

@@ 4 Diffusion model’s computation is not serial
The Serial Scaling Hypothesis provide a lens to look at machine learning from serial–parallel lens. We have examined serial machine learning problems. We now turn to a diffusion models (Ho et al., 2020; Song et al., 2021; Song and Ermon, 2019), widely used in image/video generation (Brooks et al., 2024) and other vision tasks such as depth estimation (Saxena et al., 2023), and language modeling tasks (Li et al., 2022; Nie et al., 2025; Arriola et al., 2025). While commonly thought as a serial model, we show here for the first time that they are not.
@@
While the number of denoising steps $T$ is scalable—more denoising steps for finer approximation of $p_{\text{truth}}$, in practice, diffusion models converge rapidly (Ma et al., 2025). For instance, image generation plateaus at 300 steps (Nichol and Dhariwal, 2021), depth estimation shows little difference between 5 and 100 steps (Ravishankar et al., 2024), and language modeling yields similar perplexity for 32 vs. 1024 steps (Austin et al., 2021). With distillation, 1–4 steps suffice without much loss (Yin et al., 2024; Liu et al., 2024; Song et al., 2023; Lin et al., 2024; Salimans and Ho, 2022). Such rapid convergence suggests that the effective computation depth of diffusion models is low. It would be surprising if the underlying computation were truly serial.
Diffusion models with a $\mathsf{TC}^{0}$ backbone can only solve problems in $\mathsf{TC}^{0}$. Previous work showed that a backward diffusion process converges to $p_{0}$ at the rate $TV=\mathcal{O}(d/T)$ (Li and Yan, 2024, Thm. 1), where $TV$ is the total variation between $p_{0}$ and the denoising $p_{\theta,0}$, and $d$ is the intrinsic dimension of $x_{N+1}$. Building on top of it, we obtained the following theorem, the formal statement and proof of which are in Appendix F:
@@
What does this mean? Diffusion models only provide a constant amount of additional serial computation. The above theorem precludes the use of diffusion models as a scalable means of increasing serial computation. Unlike CoT, which genuinely adds serial compute (see Appendix˜E), diffusion models do not. This may explain the empirical mediocre performance of diffusion language modeling as output length increases (Austin et al., 2021; Lou et al., 2024; Gulrajani and Hashimoto, 2023; Sahoo et al., 2024).
Limitation. The theorem doesn’t apply: (1) If solution space grows in intrinsic dimension. (2) If the backbone is either poorly trained or trained under a different objective. This leaves open the possibility of serial-compute-scalable generative models beyond score-based diffusion.
@@ 5 Acknowledgements
We thank Alexei Efros, whose critiques of deep learning theory helped ground us and motivated us to make this paper friendlier to non-theoreticians. We are grateful to William Merrill and Anant Sahai for their invaluable technical feedback and corrections that significantly strengthened our theoretical analysis. We also thank Angjoo Kanazawa, Justin Kerr, Amil Dravid, Yossi Gandelsman, and Yuatyong Chaichana for their thoughtful comments and suggestions that helped improve the clarity and presentation of this work. This work is partially supported by the ONR MURI N00014-21-1-2801. YL was supported by a gift to the Center for Human-Compatible AI at Berkeley from Coefficient Giving (formerly Open Philanthropy) during the development of this work.
@@ Appendix C Potential Misconceptions
One may be confused by the statement that “Transformers are not serial,” followed by “Chain of Thought is serial.” This appears to be a contradiction. It is not, though it may appear so due to our omission of words. The first statement really should be “If you fix a Transformer’s parameters and give it problems of size $\mathcal{O}(n)$, but only run it for $\mathcal{O}(1)$ forward passes, then you cannot solve inherently serial problem-families.”. Similarly, the second statement really should be “If you fix a Transformer’s parameters and give it problems of size $\mathcal{O}(n)$, and also run it for $\mathcal{O}(n)$ forward passes, then you can solve inherently serial problem-families.”.
This is a special case of a general phenomenon: people confuse architecture and inference methods, meshing them all together like two lumps of clay gently melding under the desert sun. This is unfortunate but understandable because, in practice, architecture and inference couple together strongly, and this confusion is also encouraged by things like “architecture–deployment co-design”. The fact is that the abstractions leak. How you use a model for inference depends on its architecture. How you design a new model architecture depends on what you expect its inference method to be. This is a good thing, but it may cause people to confuse architecture and inference.
Fundamentally, we consider an “architecture” to be a class that can have instances. For example, gpt2 is an instance of the class CausalTransformer. Each class-instance should be imagined as an “atom of computation”. An atom has mass and volume. An atom can be copied and put into many molecules. What an atom cannot do is be divided or changed. Similarly, an instance of the “Transformer architecture” is an “atomic” Transformer. This atomic Transformer does have “subatomic structures”, such as its individual layers. However, since we are not performing complicated model surgeries, such as taking off its head, reading out the tensor from its middle layers, etc., we can ignore all of its subatomic structures and assume it is simply atomic, perfectly indivisible, unchangeable, eternal. If I give you an atomic Transformer, and you want to do something with it, you must perform exactly one single forward pass, because you cannot stop half-way (since you don’t have access to its subatomic structures). This has a fixed seriality. It may be 48 layers or 96 layers, and you can pick bigger and bigger atoms from the same class with no upper bound. But for a given particular instance from the class, the seriality is $\mathcal{O}(1)$, and cannot possibly grow like $\mathcal{O}(n)$.
@@ C.2 SSM is not RNN
We class SSMs as not inherently serial, while RNNs as inherently serial. This is not a mistake, and in fact, a very valuable illustration of how an architecture can have an illusion of seriality.
@@
So they removed recurrence and preserved attention. Their title meant “attention is all you need (and recurrence is not)”. This is exactly what our paper is arguing against: attention is not all you need, and inherently serial processing is necessary. It can be done in many ways: recurrence in architecture, CoT in inference strategy, or some other method, but it cannot be avoided.
@@
Concretely, consider the problem of ingesting an input sequence: $x_{1},x_{2},\dots,x_{n}$, and the task is to output the next token $x_{n+1}$. For a Transformer with 96 layers to accomplish the task, it does so by 96 sequential steps. Each step is very wide but not deep. In contrast, an RNN with 96 layers must process $x_{1}$ first, then $x_{2}$, and so on. It requires $96n$ sequential steps that cannot be done in parallel. Suppose one attempts to do them “in parallel”, then there is an immediate problem: To run the RNN on $x_{4}$, one needs the internal states of the RNN just after it has processed $x_{3}$, and to do that, it needs to have processed $x_{2}$, etc. Abstractly, the operation of an RNN looks like:
@@
In short, because the internal states of an RNN change nonlinearly and unpredictably, one cannot skip the steps. This allows RNN to solve inherently serial problems without chain of thought, but at the price of taking wall-clock time $\mathcal{O}(n)$ to even output its first token.
@@ Appendix D $\mathsf{TC}^{0}$ and $\mathsf{TC}$ Classes
Appendix E Serial capabilities of modern machine learning methods
Empirically, we see a tradeoff in modern machine learning architectures. Some have highly parallelizable computation graphs, such as MLPs, Transformers, and SSMs, but do not solve inherently serial problems. Others can solve inherently serial problems, but cannot be parallelized, such as RNNs, repeating layers, and CoT.
@@ Appendix E Serial capabilities of modern machine learning methods
In terms of computational complexity, MLPs have been formalized as a family of Boolean circuits with threshold gates called $\sf{TC}^{0}$ (Parberry and Schnitger, 1988). Consider a family of MLPs with constant $\mathcal{O}(1)$ depth, $\mathsf{poly}(n)$ neurons per layer, and $\mathcal{O}(1)$ numerical precision. A problem is in $\sf{TC}^{0}$ if and only if it is decidable by one forward pass through an MLP under this setup.
@@
It is widely suspected that $\sf{TC}^{1}$ is strictly larger than $\sf{TC}^{0}$. Therefore, any $\sf{TC}^{1}$-hard problem requires computations deepening at a rate of at least $\mathcal{O}(\log n)$. We may consider the problem as being serial in a weaker, logarithmic sense. A simple example of $\sf{TC}^{1}$ is the word problem of the symmetric group on 5 elements $S_{5}$: Given $g_{1},g_{2},\dots,g_{n}\in S_{5}$, find $\prod_{i}g_{i}$ (Liu et al., 2022). Intuitively, this is because $S_{5}$ is not a solvable group, and therefore there is no better way to multiply its elements than via a binary tree, which has $\mathcal{O}(\log n)$ layers. Attempting to perform this with just $\mathcal{O}(1)$ layers would effectively require one to memorize the entire $S_{5}^{n}\to S_{n}$ multiplication table, which scales exponentially as $e^{n\log(120)}$, an exponential depth-width trade-off.
@@
To solve the word problem of $S_{5}$ by a Transformer in one forward pass, we present it with input tokens $g_{1},g_{2},\dots,g_{n},y$, where $y$ is a special token. The Transformer’s output at $y$ is read out as the answer. Similarly, we may solve such a problem by presenting the same input tokens to a State Space Model (SSM) or a Recurrent Neural Network (RNN). Since a Transformer with $\mathcal{O}(1)$ layers and $\mathsf{poly}(n)$ dimensions can only solve problems in $\sf{TC}^{0}$ (Merrill and Sabharwal, 2023b) in one forward pass, it cannot solve the word problem of $S_{5}$ that lies outside $\sf{TC}^{0}$.
@@
On the contrary, for an RNN, we can write the multiplication table of $S_{5}^{2}\to S_{5}$ directly into its weights, so it can solve the problem by unrolling for $\mathcal{O}(n)$ recurrence steps with $\mathcal{O}(1)$ layers and $\mathcal{O}(1)$ dimensions. Intuitively, the hidden states of an RNN keep track of the progress of multiplication as it performs the forward passes. However, RNN’s recurrence state dependency renders it non-parallelizable. Several families of SSMs were developed as a compromise that still have recurrence on hidden states, like an RNN, while making forward passes parallelizable, like a Transformer. The prototypical SSM architecture is Mamba (Gu and Dao, 2023), though there are many others.
@@
Unfortunately, it has been proven that there is not yet a “free serial-compute lunch,” in the sense that the main families of SSMs proposed so far still cannot solve the word problem of $S_{5}$ under the constraints of $\mathcal{O}(1)$ layers, $\mathsf{poly}(n)$ dimensions per layer, and one forward pass. Despite its apparent recurrence, the hidden state offered by an SSM is weaker in a computational sense than that offered by an RNN (Strobl et al., 2023; Merrill and Sabharwal, 2023a; Merrill et al., 2024). This theoretical fact rhymes with the empirical fact that in practice, MLPs, Transformers, and SSMs are all more parallelizable than RNNs.
@@
Only a genuinely serial method has been shown to go beyond the $\mathsf{TC}^{0}$ class. In addition to RNNs, this includes repeating layers and CoT. Repeating layers for $\mathcal{O}(\log n)$ times enables a standard Transformer to solve tasks in the $\mathsf{TC}^{1}$ class (Merrill and Sabharwal, 2025). With $\mathsf{poly}(n)$ CoT, requiring multiple forward passes before producing a final answer, a Transformer can be lifted from $\mathsf{TC}^{0}$ to $\mathsf{P}$ (Feng et al., 2023; Li et al., 2024). As discussed in Section˜3.5, the power of CoT has been well-attested in practice by the improved performance of reasoning models in complex math and science tasks.
@@
Appendix F Diffusion is in $\mathsf{TC}^{0}$
@@
F.1 Problem settings
An abstract language is simply a set of sentences made of letters. Formally:
@@ F.1 Problem settings
  • A sentence in an alphabet $\Sigma$ is a finite sequence of elements of $\Sigma$.
  • A language in an alphabet $\Sigma$ is a set of sentences in the alphabet $\Sigma$.
An abstract language is more than a natural language. A sentence in a natural language is a sequence of tokens. An image, divided into patches and tokenized, becomes a sentence in an abstract language. A video, divided into frames and patches and tokenized, becomes a sentence in an abstract language.
@@
We define the deterministic prefix language modeling problem: given a sequence of tokens $x_{1},\dots,x_{n}$, compute the next token $x_{n+1}$. This is a deterministic formalization of next-token prediction, the dominant paradigm in language modeling since the GPT-2 of 2019. This can also be cast into a decision problem: given input of size $n$, $x_{1},\dots,x_{n}$, decide whether $x_{n+1}$ is the correct continuation.
Most language model benchmarks can be cast into this form, at least when without chain of thought. For example, consider a problem from the GSM8K benchmark “Tina buys three 12-packs of soda for a party … How many sodas are left over when the party is over?” has only one correct continuation, “11”.
The definition has some issues. Some problems may have more than one correct answer. Some answers may occupy more than one token. We may allow chain of thought. Language modeling is not restricted to prefix modeling. A touted benefit of diffusion language modeling is that, unlike GPT models, it can generate any number of tokens anywhere in a sequence conditional on other parts of the sequence.
Define nondeterministic masked sequence modeling problem as follows. Consider a sequence of tokens $x_{1},\dots,x_{n}$, some of which are masked. The task is to compute a sequence of tokens that can acceptably fill in the masks.
For example, a game of 24 may be cast into a nondeterministic masked modeling problem [^5] Brackets would create different lengths for the masked sequence. We avoid them by reverse Polish notation. as follows: “1, 3, 9, 9, [M][M][M][M][M][M][M] = 24”. This has more than one acceptable answer, such as “9 3 ÷ 11 × 9 -”.
As another example, solving a problem with chain of thought is unmasking an entire block of masked tokens, where any unmasking is acceptable as long as the tokens between “Final answer: … [END]” are correct.
Given a problem instance, we can train a model such that, conditional on the problem instance, it would sample one point from possibly many points that form a solution manifold, from which the answer can be recovered by discretizing.
We need to make the discretization step as simple as possible, to avoid secretly performing part of the problem-solving in the discretization step. Consequently, we will only consider cases where the masking pattern is fixed, and the discretization is performed using a fixed amount of compute.
As a concrete illustration, suppose that we use a diffusion video model to evaluate an arithmetic expression. The input is a sequence of frames presenting the arithmetic expression to be evaluated, and the masked tokens are the frames illustrating an animated process that draws out the answer. The solution might be written in many variable ways, creating the solution manifold. The discretization step can be done by taking the pixel-wise nearest neighbor of the last frame’s digits.
Our overall framework for sequence modeling is:
true answersolution manifolddiscretizationstochastic sampling
We will show that if the dimension of the solution manifold does not grow quickly, and the stochastic sampling is sufficiently parallelizable and accurate, then we can derandomize it to obtain a parallelizable deterministic algorithm for obtaining the true answer, thus showing that the original problem is parallelizable, not serial.
F.2 Diffusion modeling
There are several equivalent formulations of diffusion modeling. We use the score-matching formulation. In this formulation, given a probability density $\rho_{0}$ to be modeled, and a noise schedule $\sigma_{1},\sigma_{2},\dots,\sigma_{T}$, we define the forward diffusion process by adding an increasing amount of Gaussian noise to $\rho_{0}$:
$x_{0}\sim\rho_{0},\quad x_{t}|x_{0}\sim\mathcal{N}\left(\sqrt{1-\sigma_{t}^{2}}x_{0},\sigma_{t}^{2}I\right)$
Let $\rho_{t}$ be the probability density of $x_{t}$. Any (true) score function at time $t$ is $f^{*}(t,x):=\nabla\log\rho_{t}(x)$. To score-match is to find an (approximate) score function $f_{\theta}$, such that $f_{\theta}(t,x)\approx f^{*}(t,x)$. In the sense that the average score modeling error is low:
$\epsilon_{\text{score}}:=\sqrt{\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}_{x_{t}\sim\rho_{t}}\left[\left\|f_{\theta}(t,x_{t})-f^{\star}(t,x_{t})\right\|_{2}^{2}\right]}$
Given a score approximator $f_{\theta}$ and a noise schedule $\sigma_{t}$, the Score-Matching with Langevin Dynamics (SMLD) (Song and Ermon, 2019) sampler performs a backward diffusion accordingly, sampling a sequence of $\hat{x}_{T},\hat{x}_{T-1},\dots,\hat{x}_{0}$. Let $\rho_{SMLD,t}$ be the probability density of $\hat{x}_{t}$. The theoretical basis for diffusion modeling is that, at the limit of continuous noise schedule – infinitely many steps, each adding an infinitesimal amount of noise – and the limit of perfect score function, $\rho_{SMLD,0}$ would converge to be exactly equal to $\rho_{0}$.
Generally, the difference between $\rho_{SMLD,0}$ and $\rho_{0}$ can be understood as consisting of two parts: a discretization error, caused by taking finitely many small steps, instead of infinitely many infinitesimal steps; a score-matching error part, caused by $f_{\theta}\neq f^{*}$.
Intuitively, the discretization error converges to zero as $T\to\infty$, while the score-matching error grows with $\epsilon_{\text{score}}$ and $T$. In the next section, we quote a theorem from the literature that makes this intuition precise.
F.3 The fundamental theorem (Li and Yan, 2024, Thm. 2)
Intuitively, a point is easier to model than a line, which is easier than a plane, and so on. This can be made precise by the intrinsic dimension of the support of a probability distribution to be modeled.
Definition. The support of a probability distribution $\rho_{0}$ is $\operatorname{supp}(\rho_{0})$. It is the smallest set that satisfies $\Pr_{x\sim\rho_{0}}(x\in\operatorname{supp}(\rho_{0}))=1$.
Let $T$ denote the number of SMLD denoising steps. It is a positive integer.
Definition. Given a base space $\Omega$, and two probability densities $p,q$ over it, their total variation is
$TV(p,q):=\int_{\Omega}|p(x)-q(x)|dx$
Definition. Let $X$ be an arbitrary subset of $\mathbb{R}^{n}$. Let $\epsilon>0$. The discrete intrinsic dimension is
$d_{\epsilon}(X):=\frac{\log N_{\epsilon}(X)}{\log\frac{1}{\epsilon}},$
where $N_{\epsilon}(X)$ is the minimal number of radius-$\epsilon$ balls necessary to entirely cover $X$. As $\epsilon\to 0$, if the discrete intrinsic dimension converges, then what it converges to is the intrinsic dimension.
What does it mean?
Consider a 2-dimensional square $X$ in $\mathbb{R}^{n}$. As we repeatedly halve the value of $\epsilon$, each halving would require 4 times as many radius-$\epsilon$ balls to cover the same square. Thus, $d(X)=\frac{\log 4}{\log\frac{1}{1/2}}=2$. This is true no matter if $n=3$ or $n=300$, and no matter what side length the square has. Thus, the intrinsic dimension recaptures our intuition. However, it is well-defined for more than smooth manifolds. It is well-defined for more general sets, such as fractals. The expression for intrinsic dimension is deeply related to metric entropy, and has applications in statistics and information theory. (Wainwright, 2019, Chap. 5)
However, the intrinsic dimension is not accessible in practice. For example, consider a line segment $X$ in $\mathbb{R}^{100}$. It has $d_{\epsilon}(X)\to 1$ as $\epsilon$ decreases. However, when $\epsilon$ gets small enough, suddenly $d_{\epsilon}(X)$ starts converging to 2 instead. What happened? It turns out the line segment is not really a line segment. At a high enough zooming level, it is actually a cylindrical tube. However, when $\epsilon$ got even smaller, suddenly $d_{\epsilon}(X)$ starts converging to 1 again. It turns out the cylindrical tube is actually a tightly-wound helix curve.
What is accessible in practice is the discrete intrinsic dimension $d_{\epsilon}(X)$. Intuitively, it is intrinsic dimension $X$ if we are not allowed to look more closely than $\epsilon$. Another way to intuit it is by taking $X$ and constructing a low-dimensional $\epsilon$-skeleton $X_{\epsilon}$ of it. Any point in $X$ is within distance $\epsilon$ of $X_{\epsilon}$. The intrinsic dimension of the “dehydrated skeleton” of $X$ is approximately $d_{\epsilon}(X)$.
Theorem F.1.
There exists constants $c_{M},c_{\epsilon},c>0$, such that the following is true. Fix any positive integer $T>0$. Fix $\epsilon=T^{-c_{\epsilon}}$. Let $\rho_{0}$ be the target distribution. If the target distribution has first order moment that is bounded by a polynomial in $T$:
$\mathbb{E}_{x_{0}\sim\rho_{0}}\left[\left\|x_{0}\right\|_{2}\right]\leq T^{c_{M}}.$
and we have a score network $f_{\theta}(x,t)$ such that it achieves average score modeling error $\epsilon_{\text{score}}^{2}$, then there exists a SMLD sampler schedule that takes $T$ steps, such that the result of SMLD sampling has distribution $\rho_{SMLD,0}$, and
$\mathrm{TV}\left(\rho_{0},\rho_{SMLD,0}\right)\leq c\left(\frac{d_{\epsilon}(\operatorname{supp}(\rho_{0}))}{T}(\log T)^{3}+\epsilon_{\text{score}}\sqrt{\log T}\right)$
The two terms correspond to error caused by taking discrete steps, and error caused by drifting away from score matching error.
F.4 Main theorem
We formalize a model of non-uniform $\mathsf{TC}^{0}$ diffusion modeling.
Given a task, we define what it means to solve it by nondeterministic masked sequence modeling. Each task instance is specified by a sequence of $n$ tokens, written as $x_{1},\dots,x_{n}$. Of these, $k_{n}$ tokens are fixed inputs to the model to specify the problem, and $n-k_{n}$ tokens are masked, which the model would denoise. The denoised tokens are discretized to a final answer.
To avoid surreptitiously hinting the diffusion model the answer, the masked tokens are always in the same place, which, WLOG, we assume comes at the end. Similarly, we require that for each $n$, $k_{n}$ is fixed. Finally, we require the discretization step to require constant compute, to avoid secretly offloading problem-solving computation from the diffusion model to the discretization algorithm.
Given a task, a $\mathsf{TC}^{0}$ family of score-networks $f_{\theta,n}$ for such a task is a sequence of networks $f_{\theta,1},f_{\theta,2},\dots$, such that: